solve
James has saved $560 in two years. He saved 3 times as much during the first year as during the second year. How much did he save each year?

Answer by Paul(988)

Let the second year be x
let the first year be 3x
sum = 560
3x+x=560
4x=560
4x/4=560/4
x=140
3x140= 420
Therefore he saved 420 first year and 140 second year